# Four-dimensional gradient Ricci solitons with (half) nonnegative isotropic curvature

### v2

@@ 2.3. Some basic facts about Ricci solitons and ancient Ricci flows
Thus, by Hamilton’s strong maximum principle (see, e.g., [25, Theorem 2.2.1]), there exists an interval $0<t<\delta$ over which the rank of $A(t)$ is constant, and $\ker(A(t))$ is invariant under parallel translation and invariant in time. As $g(t)=\Phi(t)^{\ast}(g)$ with $g(0)=g$, we conclude that $\ker(A)=\ker(A(0))$ is also invariant under parallel translation for $(M^{n},g,f)$. Similarly, $\ker(C)$ is invariant under parallel translation.
Next, we show that if $(M^{4},g,f)$ has WPIC and the holonomy group $\text{Hol}^{0}(M^{4},g)$ is $\text{SO}(4)$, then $A>0$ and $C>0$ hence $(M^{4},g,f)$ has PIC and also $Rc>0$ by Theorem 3.1(b). Actually, this follows from [8, Corollary 9.14]. However, for the sake of the proof of Theorem 1.3, we provide another argument below that also works for the half WPIC case.
We essentially follow the same argument as in the proof of [24, Theorem 1.2] and argue by contradiction. Suppose that there exist a point $p\in M^{4}$ and a self-dual $\varphi_{1}\in\wedge^{+}(M_{p})$ such that $A(\varphi_{1},\varphi_{1})=0.$ It is then clear that $\varphi_{1}$ is a null eigenvector corresponding to the smallest eigenvalue $A_{1}=0$ at $p$. Now, by [44, Lemma 6.1], the self-dual $\varphi_{1}\in\wedge^{+}(M_{p})$ can be expressed as
@@
for some positively oriented orthonormal frame $\{e_{1},e_{2},e_{3},e_{4}\}$. Meanwhile, suppose $\varphi_{3}\in\wedge^{+}(M_{p})$ is an eigenvector corresponding to the largest eigenvalue $A_{3}$ at $p$. Then, using [44, Lemma 6.1] again, we can find another positively oriented orthonormal frame $\{v_{1},v_{2},v_{3},v_{4}\}$ such that
@@
where $P_{\gamma}$ denotes the parallel transport along $\gamma$. It then follows that $A_{3}=A(\varphi_{3},\varphi_{3})$ $=A(\varphi_{1},\varphi_{1})=A_{1}=0$, since $\ker(A)$ is invariant under parallel translation. This would imply that the scalar curvature $R=4(A_{1}+A_{2}+A_{3})=0$ at $p$. By Remark 2.2, $(M^{4},g,f)$ must be Ricci flat. However, together with the third inequality in Lemma 2.1, this would imply $(M^{4},g,f)$ is flat, a contradiction to the assumption that $\text{Hol}^{0}(M^{4},g)=\text{SO}(4)$. Thus, $A>0$. Similarly, $C>0$. Therefore, $(M^{4},g,f)$ has PIC and hence also $Rc>0$ by Theorem 3.1(b).
Now, we are ready to conclude the proof.
Case 2: $(M^{4},g,f)$ is locally irreducible and locally symmetric. Then, as $(M^{4},g,f)$ is a steady Ricci soliton it must be Ricci flat and hence flat due to WPIC as above, or by [5, Theorem 7.61]. So, Case 2 cannot occur when $(M^{4},g,f)$ is assumed non-flat.
Case 3: $(M^{4},g,f)$ is locally irreducible and not locally symmetric. Then, by Berger’s holonomy classification, either $\text{Hol}^{0}(M^{4},g)=\text{SO}(4)$, or $\text{Hol}^{0}(M^{4},g)=\text{U}(2)$, or $\text{Hol}^{0}(M^{4},g)=\text{SU}(2)$. If $\text{Hol}^{0}(M^{4},g)=\text{SO}(4)$ then, from the above, we know that $(M^{4},g,f)$ must have PIC and positive Ricci curvature. On the other hand, if $\text{Hol}^{0}(M^{4},g)=\text{SU}(2)$ then $(M^{4},g,f)$ is Calabi-Yau, hence flat due to WPIC. But this is ruled out by the non-flatness assumption. Finally, if $\text{Hol}^{0}(M^{4},g)=\text{U}(2)$, then $(M^{4},g)$ is Kähler, hence $(M^{4},g,f)$ is a gradient steady Kähler-Ricci soliton. On the other hand, any gradient steady Kähler-Ricci soliton of complex dimension two with WPIC necessarily has nonnegative curvature operator; see, e.g., [34, Lemma 4.6]. Moreover, since $(M^{4},g,f)$ is locally irreducible, not locally symmetric and has $\text{Hol}^{0}(M^{4},g)=\text{U}(2)$, by using [46, Theorem 8.3] and a similar argument as in [20], it follows that $(M,g)$ has $Rm>0$ when restricted to real $(1,1)$-forms $\wedge^{1,1}_{\mathbb{R}}(M)$.
This completes the proof of Theorem 3.2/Theorem 1.2. ∎
Proof of Theorem 1.3. First of all, by Lemma 2.7, $(M^{4},g,f)$ has either $A\geq 0$ or $C\geq 0$. Now, by essentially the same argument as in the proof of Theorem 3.2, it follows that either $(M^{4},g,f)$ has half PIC, or it is a gradient Kähler-Ricci steady soliton, or it is Ricci-flat, or it splits as a product. Indeed,
@@
Case 3: $(M^{4},g,f)$ is locally irreducible and non-symmetric. Then, by Berger’s holonomy classification, either $\text{Hol}^{0}(M^{4},g)=\text{SO}(4)$, or $\text{Hol}^{0}(M^{4},g)=\text{U}(2)$, or $\text{Hol}^{0}(M^{4},g)=\text{SU}(2)$. If $\text{Hol}^{0}(M^{4},g)=\text{SO}(4)$ then, from the proof of Theorem 3.2 above, $(M^{4},g,f)$ either is Ricci flat or has half PIC. If it is Ricci flat, then it implies that $A=W^{+}$ and $C=W^{-}$ because scalar curvature $R=0$. However, by Lemma 2.7, we also have either $A\geq 0$ or $C\geq 0$. Moreover, $\text{tr}W^{+}=\text{tr}W^{-}=0$. Hence, it follow that either $W^{+}=0$ or $W^{-}=0$, i.e., $(M^{4},g,f)$ is either anti-self-dual (ASD) or self-dual (SD). Moreover, it is well-known that Ricci-flat and ASD imply that the universal cover is hyperKähler. On the other hand, if $\text{Hol}^{0}(M^{4},g)=\text{U}(2)$, then $(M^{4},g,f)$ is a gradient steady Kähler-Ricci soliton. Finally, if $\text{Hol}^{0}(M^{4},g)=\text{SU}(2)$ then $(M^{4},g,f)$ is Calabi-Yau, and its universal cover is hyperKähler.
This finishes the proof of Theorem 1.3. $\square$
@@
Next, we claim that if the holonomy group $\text{Hol}^{0}(M^{4},g)$ is $\text{SO}(4)$, then $(M^{4},g,f)$ has PIC. Indeed, consider the canonical Ricci flow $g(t)=(1+t)\Phi(t)^{\ast}(g)$ induced by the expanding soliton $(M^{n},g,f)$ with $g(0)=g$ for $t\in[0,1]$. By [12, Proposition 8], for any fixed time $\tau\in(0,1)$, the set of all orthonormal four-frames $\{e_{1},e_{2},e_{3},e_{4}\}$ on $(M^{4},g(\tau))$ satisfy
@@
is invariant under parallel transport with respect to the metric $g(\tau)$. Moreover, as $g(t)=(1+t)\Phi(t)^{\ast}(g)$ and $g(0)=g$, we conclude that all orthonormal four-frames $\{e_{1},e_{2},e_{3},e_{4}\}$ on $(M^{n},g,f)$ satisfying (4.3) is also invariant under parallel translation with respect to $g$. Thus, by [8, Corollary 9.14], if the holonomy group $\text{Hol}^{0}(M^{4},g)$ is $\text{SO}(4)$ then $(M^{4},g,f)$ has PIC.
@@
This finishes the proof of Theorem 4.1/Theorem 1.4. ∎
@@ 5. The improved results for 4D shrinking Ricci solitons
Let $(M^{4},g,f)$ be an oriented 4-dimensional complete gradient shrinking Ricci soliton with $A_{2}\geq 0$ on $M^{4}$. Then, $(M^{4},g,f)$ either has $A_{2}>0$ on $M^{4}$ (hence $A>0$), or is isometric to the Gaussian soliton $\mathbb{R}^{4}$ or a $\mathbb{Z}_{2}$ quotient of $\mathbb{S}^{2}\times\mathbb{S}^{2}$ or $\mathbb{S}^{2}\times\mathbb{R}^{2}$, or is a gradient shrinking Kähler-Ricci soliton.
@@
For oriented positive Einstein $4$-manifold $(M^{4},g)$, recently Li and Zhang [60] showed that if $H^{1}(M,\mathbb{Z}_{2})=0$ and $(M^{4},g)$ is not isometric to a round 4-sphere $\mathbb{S}^{4}$, then $(M^{4},g)$ is Kähler-Einstein if and only if $A_{2}\geq 0$ on $M^{4}$; furthermore, if $A_{2}>0$ and $M^{4}$ is simply connected then $(M^{4},g)$ is isometric to $\mathbb{S}^{4}$ with the round metric. The proof in [60] is based on the Weitzenböck formula from Micallef-Wang [64] for $W^{+}$ with $\delta W^{+}=0$. Note that the combination of our Proposition 5.1(b) and [64, Theorem 4.2] also gives an alternative proof.
@@ 6. Further Discussions
As we have seen in Section 3, Lemma 2.7 played an essential role in the proof of Theorem 3.1. On the other hand, by using the WPIC condition (i.e., $A$ and $C$ are weakly 2-positive) and a similar argument as in the proof of our Theorem 3.1(a), it turns out that the property of $Rc$ being 2-nonnegative can be shown to be preserved by 4D Ricci flow with WPIC.
@@
By (2.4) and Lemma 2.2, the sum of the two least eigenvalues of $Rc$ is given by
$\Lambda_{1}+\Lambda_{2}=-2B_{3}+\frac{R}{2}.$
Thus, 2-nonnegative Ricci curvature is equivalent to $R-4B_{3}\geq 0.$
Now, we compute the ODE of $R-4B_{3}$. By Lemma 2.4 and $|\mathring{Rc}|^{2}=4|B|^{2}$, we have
$\begin{split}\frac{d}{dt}\left(R-4B_{3}\right)&\geq 2|Rc|^{2}-8(A_{3}B_{3}+C_{3}B_{3}+2B_{1}B_{2})\\ &=\frac{1}{2}R^{2}+8|B|^{2}-8(A_{3}B_{3}+C_{3}B_{3}+2B_{1}B_{2})\\ &=\frac{1}{2}\left(R-4B_{3}\right)^{2}+4RB_{3}+8(B_{2}-B_{1})^{2}-8(A_{3}+C_{3})B_{3}\\ &\geq\frac{1}{2}\left(R-4B_{3}\right)^{2}+4(R-2A_{3}-2C_{3})B_{3}.\end{split}$
On the other hand, since $A_{1}+A_{2}+A_{3}=C_{1}+C_{2}+C_{3}=\frac{R}{4}$, we have
$R-4A_{3}=4(A_{1}+A_{2})\qquad\mbox{and}\qquad R-4C_{3}=4(C_{1}+C_{2}).$
Hence, we obtain
@@
where, in the second inequality, we have used the assumption of WPIC, i.e., $A_{1}+A_{2}\geq 0$ and $C_{1}+C_{2}\geq 0$. Therefore, by Hamilton’s maximum principle [46, Theorem 4.2], the 2-nonnegative Ricci curvature condition is preserved under the 4-dimensional Ricci flow with WPIC.
This finishes the proof of Proposition 6.1. ∎
@@
As we mentioned before, for $n\geq 5$, Li and Ni [58] showed that complete ancient solutions to the Ricci flow with WPIC must have 2-nonnegative Ricci curvature; see [58, Proposition 5.2]. By using essentially the same argument as in their proof, it follows that the 2-nonnegativity of $Rc$ is preserved by $n$-dimensional $(n\geq 5$) Ricci flow with WPIC too. However, their proof does not seem to extend to the $n=4$ case, since the key inequality in [58, Lemma 4.1] holds only for $n\geq 5$.
Next, we observe that the condition of 2-nonnegative Ricci curvature is preserved by the Ricci flow in dimension three. In this case, we can diagonalize the curvature operator $Rm$ with eigenvalues $m_{1}\leq m_{2}\leq m_{3}$. Then the Ricci tensor $Rc$ is diagonalized with eigenvalues $\frac{1}{2}(m_{1}+m_{2})\leq\frac{1}{2}(m_{1}+m_{3})\leq\frac{1}{2}(m_{2}+m_{3})$, and the scalar curvature is given by $R=m_{1}+m_{2}+m_{3}$. Moreover, we have the following ODE system corresponding to the curvature evolution PDE ([46]):
@@
Therefore, by Hamilton’s maximum principle [46, Theorem 4.2], the 2-nonnegative Ricci curvature condition is preserved by the 3-dimensional Ricci flow. ∎
@@
for some constant $\Lambda>0$. Furthermore, they showed that such ancient solutions necessarily have bounded curvature, provided they are $\kappa$-noncollapsed. Combining these properties with the result [9, Corollary 1.6] of Brendle and Naff, it follows that any $4$-dimensional $\kappa$-noncollapsed, noncompact, complete ancient solution to the Ricci flow with uniformly PIC is isometric to either a family of shrinking cylinders (or their quotients) or the Bryant soliton.
@@
On the other hand, motivated by the result of B.-L. Chen [31] that any $3$-dimensional complete ancient solution to the Ricci flow must have $Rm\geq 0$, it is also natural to make the following
Conjecture/Question 6.1555Indeed, this question was raised by the first author at a colloquium talk given at Sun Yat-sen University (Guangzhou, China) in December, 2018.. Let $g(t)$ be a $4$-dimensional complete ancient solution to the Ricci flow on $M^{4}$ with PIC (or WPIC). Is it true that $g(t)$ necessarily has either nonnegative curvature operator $Rm\geq 0$ or WPIC2?
@@
(6.3) $A_{3}\leq L(A_{1}+A_{2})\quad{\mbox{and}}\quad C_{3}\leq L(C_{1}+C_{2}),$

### v3

@@ 2.3. Some basic facts about Ricci solitons and ancient Ricci flows
Thus, by Hamilton’s strong maximum principle (see, e.g., [25, Theorem 2.2.1]), there exists an interval $0<t<\delta$ over which the rank of $A(t)$ is constant, and $\ker(A(t))$ is invariant under parallel translation and invariant in time. As $g(t)=\Phi(t)^{\ast}(g)$ with $g(0)=g$, we conclude that $\ker(A)=\ker(A(0))$ is also invariant under parallel translation for $(M^{n},g,f)$.
Claim: If $(M^{4},g,f)$ has half WPIC and the restricted holonomy group $\text{Hol}^{0}(M^{4},g)$ is $\text{SO}(4)$, then either it is Ricci flat, or $A>0$ hence $(M^{4},g,f)$ has half PIC.
We essentially follow the same argument as in the proof of [24, Theorem 1.2] and argue by contradiction. Suppose that $(M^{4},g,f)$ is not Ricci flat and there exist a point $p\in M^{4}$ and a self-dual $\varphi_{1}\in\wedge^{+}(M_{p})$ such that $A(\varphi_{1},\varphi_{1})=0.$ It is then clear that $\varphi_{1}$ is a null eigenvector corresponding to the smallest eigenvalue $A_{1}=0$ at $p$. Now, by [43, Lemma 6.1], the self-dual $\varphi_{1}\in\wedge^{+}(M_{p})$ can be expressed as
@@
for some positively oriented orthonormal frame $\{e_{1},e_{2},e_{3},e_{4}\}$. Meanwhile, suppose $\varphi_{3}\in\wedge^{+}(M_{p})$ is an eigenvector corresponding to the largest eigenvalue $A_{3}$ at $p$. Then, using [43, Lemma 6.1] again, we can find another positively oriented orthonormal frame $\{v_{1},v_{2},v_{3},v_{4}\}$ such that
@@
where $P_{\gamma}$ denotes the parallel transport along $\gamma$. It then follows that
$A_{3}=A(\varphi_{3},\varphi_{3})=A(\varphi_{1},\varphi_{1})=A_{1}=0,$
since $\ker(A)$ is invariant under parallel translation. This would imply that the scalar curvature $R=4(A_{1}+A_{2}+A_{3})=0$ at $p$. By Remark 2.2, $(M^{4},g,f)$ must be Ricci flat, which is a contradiction. Thus, $A>0$. Therefore, $(M^{4},g,f)$ has PIC and the Claim is proved.
Now, we are ready to finish the proof of Theorem 3.2
@@
Case 3: $(M^{4},g,f)$ is locally irreducible and non-symmetric. Then, by Berger’s holonomy classification, either $\text{Hol}^{0}(M^{4},g)=\text{SO}(4)$, or $\text{Hol}^{0}(M^{4},g)=\text{U}(2)$, or $\text{Hol}^{0}(M^{4},g)=\text{SU}(2)$. If $\text{Hol}^{0}(M^{4},g)=\text{SO}(4)$ then, from the Claim above, $(M^{4},g,f)$ either is Ricci flat or has half PIC. If it is Ricci flat, then it implies that $A=W^{+}$ and $C=W^{-}$ because scalar curvature $R=0$. However, by Lemma 2.7, we also have either $A\geq 0$ or $C\geq 0$. Moreover, $\text{tr}W^{+}=\text{tr}W^{-}=0$. Hence, it follow that either $W^{+}=0$ or $W^{-}=0$, i.e., $(M^{4},g,f)$ is either anti-self-dual (ASD) or self-dual (SD). Moreover, it is well-known that Ricci-flat and ASD imply that the universal cover is hyperKähler. On the other hand, if $\text{Hol}^{0}(M^{4},g)=\text{U}(2)$, then $(M^{4},g,f)$ is a gradient steady Kähler-Ricci soliton. Finally, if $\text{Hol}^{0}(M^{4},g)=\text{SU}(2)$ then $(M^{4},g,f)$ is Calabi-Yau, and its universal cover is hyperKähler.
This finishes the proof of Theorem 3.2 (Theorem 1.3). ∎
Remark 3.2.
As mentioned in the introduction, Theorem 1.2 and Theorem 1.3 are valid under some slightly weaker assumption than WPIC (or half WPIC); see Remark 5.3.
@@
Next, we claim that if the holonomy group $\text{Hol}^{0}(M^{4},g)$ is $\text{SO}(4)$, then $(M^{4},g,f)$ has PIC. Indeed, consider the canonical Ricci flow $g(t)=(1+t)\Phi(t)^{\ast}(g)$ induced by the expanding soliton $(M^{n},g,f)$ with $g(0)=g$ for $t\in[0,1]$. As in the proof of Theorem 1.2, by [12, Proposition 8], for any fixed time $t\in(0,1)$, the set of all orthonormal four-frames $\{e_{1},e_{2},e_{3},e_{4}\}$ on $(M^{4},g(t))$ satisfy
@@
is invariant under parallel transport with respect to the metric $g(t)$, which implies that all orthonormal four-frames $\{e_{1},e_{2},e_{3},e_{4}\}$ on $(M^{n},g,f)$ satisfying (4.3) is also invariant under parallel translation with respect to $g$. Thus, by [8, Corollary 9.14], if the holonomy group $\text{Hol}^{0}(M^{4},g)$ is $\text{SO}(4)$ then $(M^{4},g,f)$ has PIC.
@@
This finishes the proof of Theorem 4.1 (Theorem 1.4). ∎
@@ 5. The improved results for 4D shrinking Ricci solitons
Let $(M^{4},g,f)$ be an oriented 4-dimensional complete gradient shrinking Ricci soliton with $A_{2}\geq 0$ on $M^{4}$. Then, $(M^{4},g,f)$ either has $A>0$ on $M^{4}$, or is isometric to the Gaussian soliton $\mathbb{R}^{4}$ or a $\mathbb{Z}_{2}$ quotient of $\mathbb{S}^{2}\times\mathbb{S}^{2}$ or $\mathbb{S}^{2}\times\mathbb{R}^{2}$, or is a gradient shrinking Kähler-Ricci soliton.
@@
For oriented positive Einstein $4$-manifold $(M^{4},g)$, recently Li and Zhang [59] showed that if $H^{1}(M,\mathbb{Z}_{2})=0$ and $(M^{4},g)$ is not isometric to a round 4-sphere $\mathbb{S}^{4}$, then $(M^{4},g)$ is Kähler-Einstein if and only if $A_{2}\geq 0$ on $M^{4}$; furthermore, if $A_{2}>0$ and $M^{4}$ is simply connected then $(M^{4},g)$ is isometric to $\mathbb{S}^{4}$ with the round metric. The proof in [59] is based on the Weitzenböck formula from Micallef-Wang [62] for $W^{+}$ with $\delta W^{+}=0$. Note that the combination of our Proposition 5.1(b) and [62, Theorem 4.2] also gives an alternative proof.
@@ 6. Further Discussions
By using the WPIC condition (i.e., $A$ and $C$ are weakly 2-positive), the same computation as in the proof of our Theorem 3.1(a) and Hamilton’s maximum principle [45, Theorem 4.2] imply that the 2-nonnegative Ricci curvature property is preserved by 4D Ricci flow with WPIC.
@@
By Lemma 2.2, we know that 2-nonnegative Ricci curvature is equivalent to
$R-4B_{3}\geq 0.$
Moreover, as we have seen in the proof of Theorem 3.1(a),
@@
where, in the second inequality, we have used the assumption of WPIC, i.e., $A_{1}+A_{2}\geq 0$ and $C_{1}+C_{2}\geq 0$. Therefore, by Hamilton’s maximum principle [45, Theorem 4.2], the 2-nonnegative Ricci curvature condition is preserved under the 4-dimensional Ricci flow with WPIC. This finishes the proof of Proposition 6.1. ∎
@@
As we mentioned before, for $n\geq 5$, Li and Ni [57] showed that complete ancient solutions to the Ricci flow with WPIC must have 2-nonnegative Ricci curvature; see [57, Proposition 5.2]. By using essentially the same argument as in their proof, it follows that the 2-nonnegativity of $Rc$ is also preserved by $n$-dimensional $(n\geq 5$) Ricci flow with WPIC. However, their proof does not seem to extend to the $n=4$ case, since the key inequality in [57, Lemma 4.1] holds only for $n\geq 5$.
Next, we observe that the condition of 2-nonnegative Ricci curvature is preserved by the Ricci flow in dimension three. In this case, we can diagonalize the curvature operator $Rm$ with eigenvalues $m_{1}\leq m_{2}\leq m_{3}$. Then the Ricci tensor $Rc$ is diagonalized with eigenvalues
$\frac{1}{2}(m_{1}+m_{2})\leq\frac{1}{2}(m_{1}+m_{3})\leq\frac{1}{2}(m_{2}+m_{3}),$
and the scalar curvature is given by $R=m_{1}+m_{2}+m_{3}$. Moreover, we have the following ODE system corresponding to the curvature evolution PDE ([45]):
@@
Therefore, by Hamilton’s maximum principle [45, Theorem 4.2], the 2-nonnegative Ricci curvature condition is preserved by the 3-dimensional Ricci flow. ∎
@@
for some constant $\Lambda>0$. Furthermore, they showed that such ancient solutions necessarily have bounded curvature, provided they are $\kappa$-noncollapsed. Combining these properties with the result [11, Corollary 1.6] of Brendle and Naff, it follows that any $4$-dimensional $\kappa$-noncollapsed, noncompact, complete ancient solution to the Ricci flow with uniformly PIC is isometric to either a family of shrinking cylinders (or their quotients) or the Bryant soliton.
@@
On the other hand, motivated by the result of B.-L. Chen [30] that any $3$-dimensional complete ancient solution to the Ricci flow must have $Rm\geq 0$, it is also natural to ask the following
Question 6.1666This question was raised by the first author at a colloquium talk given at Sun Yat-sen University (Guangzhou, China) in December, 2018.. Let $g(t)$ be a $4$-dimensional complete ancient solution to the Ricci flow on $M^{4}$ with PIC (or WPIC). Is it true that $g(t)$ necessarily has nonnegative curvature operator $Rm\geq 0$?
@@
(6.3) $A_{3}\leq L(A_{1}+A_{2})\qquad{\mbox{and}}\qquad C_{3}\leq L(C_{1}+C_{2}),$
@@
Acknowledgements. We are grateful to Professor Richard Hamilton for helpful suggestions. We would also like to thank Dr. Pak-Yeung Chan, Dr. Florian Johne, Dr. Yi Lai, and the anonymous referee for helpful comments. The first author’s research was partially supported by a grant from the Simons Foundation.